# SRA-MCTS: Self-driven Reasoning Aurmentation with Monte Carlo Tree Search for Enhanced Code Generation

### v1

@@ Abstract
SRA-MCTS: Self-driven Reasoning Aurmentation with Monte Carlo Tree Search for Enhanced Code Generation
@@ 1 Introduction
Some work enhances the answers in the data by incorporating a natural language intermediary process as part of the response, guiding the model to generate results in the correct direction. Some use GPT3.5 to generate natural language solutions in a step-by-step Chain-of-Thought (CoT) format Wang et al. (2024b). Some simply introduce guiding ideas to steer the process Li et al. (2023). The previous work aims to guide the large model’s reasoning process toward generating correct results through natural language as an intermediary. However, the inherent uncertainty in the model’s generation process means that errors may still arise during the formulation of these natural language ideas.
The experiments conducted by ScaleAI provide concrete experimental validation for previous work on answers that providing LLMs with correct solutions in natural language as a part of the answer, even if incomplete (just 10-20 tokens), can substantially boost the performance on benchmarks Wang et al. (2024a). In contrast, incorrect solutions result in a decline in solve rate. This demonstrates that providing solutions to large models can guide and inspire their reasoning process, and the correctness of the solution directly impacts the accuracy of the final result.

### v3

# SRA-MCTS: Self-driven Reasoning Augmentation with Monte Carlo Tree Search for Code Generation

@@ Abstract
SRA-MCTS: Self-driven Reasoning Augmentation with Monte Carlo Tree Search for Code Generation
@@ 1 Introduction
Some work enhances the answers in the data by incorporating a natural language intermediary process as part of the response, guiding the model to generate results in the correct direction. Some use GPT3.5 to generate natural language solutions in a step-by-step Chain-of-Thought (CoT) format Wang et al. (2024b). Some simply introduce guiding ideas to steer the process Li et al. (2023). The previous work aims to guide the LLMs’ reasoning process toward generating correct results through natural language as an intermediary. However, the inherent uncertainty in the model’s generation process means that errors may still arise during the formulation of these natural language ideas.
ScaleAI Wang et al. (2024a) provides concrete experimental validation for previous work on answers that providing LLMs with correct solutions in natural language as a part of the answer, even if incomplete (just 10-20 tokens), can substantially boost the performance on benchmarks. In contrast, incorrect solutions result in a decline in solve rate. This demonstrates that providing solutions to LLMs can guide and inspire their reasoning process, and the correctness of the solution directly impacts the accuracy of the final result.